Write the conic in standard form.
y^2-x^2+2y-14x-57=0

Answer by lwsshak3(11628)

Write the conic in standard form.
y^2-x^2+2y-14x-57=0
complete the square
(y^2+2y+1)-(x^2+14x+49)=57+1-49
(y+1)^2-(x+7)^2=9
(y+1)^2/9-(x+7)^2/9=1
This is an equation of a hyperbola with vertical transverse axis.
Its standard form of equation: (y-k)^2/a^2-(x-h)^2=1, (h,k)=(x,y) coordinates of center.
For given equation: (y+1)^2/9-(x+7)^2/9=1
center: (-7,-1)
a^2=9
b^2=9